OpenStudy (anonymous):

Two airfields A and B are 400 miles apart, and B is due east of A.A plane flew from A to B in two hours and then returned to A in 2 1/2 hours.If the wind blew with a constant velocity from the west during the entire trip,find the speed of the plane in still air and the speed of the wind?

OpenStudy (anonymous):

please help with solutions.

OpenStudy (anonymous):

say speed of plane=v speed of wind=v' 400/(v+v') = 5/2 hours 400/(v-v') = 2 hours solve these for v and v'